OpenStudy (anonymous):

Find the vector equation and parametric equations for L2, the straight line passing through the point A(4; 1; 2), and parallel to the line described by the symmetric equation x-3=(y-6)/3=(z-1)/3

OpenStudy (anonymous):

Not meant to be semicolons between the numbers on point A sorry. I know i need to use the equation \[R=r+tv\] but i'm not sure how to get the symmetric equation into a point where i can use it.

OpenStudy (anonymous):

x-3=(y-6)/3=(z-1)/3=t so: x-3=t y-6=3t z-1=3t or: if r=(3,6,1) and v=(1,3,3), then: R=(3,6,1)+t(1,3,3)

OpenStudy (anonymous):

Thanks for the answer. Sorry to ask but then how does point A come into the question?

OpenStudy (anonymous):

what I gave you is the vector expretion og the line x-3=(y-6)/3=(z-1)/3. It is not the final answer. You have to take the direction vector from this line and aply it to the point A to get the final answer

OpenStudy (anonymous):

Okay that makes sense. thanks.

OpenStudy (anonymous):

yw